# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition boxthm]Assumption

### v1

@@ 3.2 Fair Optimal Transport
Under Assumption 3.2, for any $\mathbf{F}\in\Pi(\mathbf{p},\mathbf{q})$, there exists a unique fair optimal transport plan.
@@ 3.3 Solving Fair Optimal Transport
where
$\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}\kern 5.0pt.$
This result mirrors the foundational insight behind the Sinkhorn algorithm and shows that the fairness constraint can be incorporated via a simple additional projection step. This leads to our modified Sinkhorn algorithm, which we call FairSinkhorn, presented in Algorithm˜1 in which the function $\Phi$ is defined as
$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
The supplementary steps with respect to the original Sinkhorn algorithm are highlighted in blue.
@@
5:  $\mathbf{u}^{(t+1)}\leftarrow\mathbf{a}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})\mathbf{v}^{(t)}\Big)$
6:  $\mathbf{v}^{(t+1)}\leftarrow\mathbf{b}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})^{\top}\mathbf{u}^{(t+1)}\Big)$
@@
8:  $\mathbf{T}^{(t+1)}\leftarrow\sum_{sw}\ell^{(t+1)}_{sw}\mathbf{B}_{sw}$ where $\ell^{(t+1)}_{sw}=\big(\mathbf{L}^{(t+1)}\big)_{sw}$
@@ 4 Two Strategies for Approximately Fair OT
While the previous section showed that exact fairness can be enforced via a projection step in FairSinkhorn, strictly satisfying fairness constraints can substantially increase transport cost, sometimes beyond what is acceptable in practice. This price of fairness has been studied in operations research for resource allocation problems [bertsimas2011price] and also arises in supervised learning, where enforcing exact group fairness constraints may require sacrificing predictive accuracy [zhao2019inherent]. Motivated by these observations, it is often preferable in practice to allow some tolerance, trading off slight deviations from fairness against overall efficiency.
In this section, we implement this idea by replacing the linear fairness constraints in (LABEL:eq:pure_empirical_fairness) with a relaxed condition
$\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}):=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\bigg(\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\bigg)^{2}\leq\rho,$ (4)
where $\rho\geq 0$ is a tolerance level. This formulation allows the transport plan to deviate slightly from the exact fairness target while still controlling the total fairness violation. To find optimal transport plans under this relaxed constraint, we propose two strategies.
@@ 4.1 Fairness-Penalized OT
where $\lambda>0$ controls the strength of the fairness penalty. Importantly, since $\mathcal{L}_{\mathbf{F}}$ is convex, the objective remains strongly convex, ensuring a unique minimizer. This penalized entropic optimal transport problem can be solved efficiently using a generalized conditional gradient algorithm [rakotomamonjy2015generalized]. At each iteration, the algorithm linearizes the fairness regularization around the current iterate and solves a subproblem over the feasible set $\Pi$ to find a search direction. A line search then determines a convex combination of the current iterate and this direction, guaranteeing descent. In practice, the subproblem at iteration $t$ corresponds to an entropic OT problem with the modified cost $\mathbf{C}+\nabla\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}^{t})$ which is solved efficiently with the Sinkhorn algorithm. This makes the algorithm computationally practical even for large-scale problems.
Sample Complexity of Fairness-Penalized OT.
An important contribution of our work is to establish a sample complexity bound for the above fairness-penalized entropic optimal transport cost, by building upon results from entropic optimal transport [genevay2019sample, mena2019statistical]. To formalize this, note that the optimization problem in (5) can be defined for arbitrary measures $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ as (see Section˜B.2 for a formal definition)
$m^{\star}(\mu,\eta):=\min_{\pi\in\Pi(\mu,\eta)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)\kern 5.0pt.$
Let $\mu_{n}=\sum_{i=1}^{n}\delta_{x_{i}}$ and $\eta_{n}=\sum_{i=1}^{n}\delta_{y_{i}}$ where $(x_{i})_{1\leq i\leq n}$ and $(y_{i})_{1\leq i\leq n}$ are i.i.d. samples drawn from $\mu$ and $\eta$. In Section˜4.1, we quantify how well we can estimate $m^{\star}(\mu,\eta)$ with $m^{\star}(\mu_{n},\eta_{n})$ in function of the sample size $n$. Similar to [rigollet2022sample] and [genevay2019sample], our proof relies on the fact that the measures are compactly supported. We also require the following assumption on the ground cost function.
@@
Compared to existing sample complexity bounds for the standard (non-penalized) entropic OT problem [genevay2019sample, mena2019statistical], our result attains the same scaling with the number of samples $n$, demonstrating that introducing the fairness penalty does not worsen the statistical efficiency of the estimator. The proof can be extended to generic convex penalties at the price of a supplementary $\log(n)$ factor.
@@
$Y_{n}-m^{\star}(\mu,\eta)\leq m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq Z_{n}-m^{\star}(\mu,\eta)\kern 5.0pt.$
Lower bound. We obtain $Y_{n}$ through linearization of the convex penalty, which is guaranteed to yield a proper lower bound through the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized]. This linearized problem may then be cast as an entropic optimal transport problem with modified cost, allowing us to leverage standard results on the sample complexity of entropic optimal transport [genevay2019entropy, mena2019statistical].
Upper bound. To obtain $Z_{n}$ we consider $\hat{\pi}_{n}^{\star}$, the minimizer of the linearized problem, and insert it in the fairness-penalized loss function, that is we consider
$m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq\langle c,\hat{\pi}_{n}^{\star}\rangle+\varepsilon\textbf{KL}(\hat{\pi}_{n}^{\star}|\mu_{n}\otimes\eta_{n})+\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta),$ (6)
where the upper bound follows from the fact that $\hat{\pi}_{n}^{\star}$ satisfies the constraints defining $m^{\star}(\mu_{n},\eta_{n})$. To establish that, in expected value, the absolute value of (6) vanishes at the rate of $1/\sqrt{n}$, we again make use of the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized] to reduce the problem to the sample complexity of entropic optimal transport. We handle the part $\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta)$ by considering a generalization of Theorem 6 from [rigollet2022sample]. ∎
@@ 4.2 Fair OT via Cost Learning
The key idea behind our second strategy is to learn a ground cost function $c_{\theta}$ so that the entropic optimal transport (OT) solution $\mathbf{\Pi}(c_{\theta})$ induced by this cost minimizes the fairness violation. In other words, rather than modifying the transport plan directly, we reshape the geometry of the OT problem itself. To control deviations from the original OT problem, we require the learned cost to remain close to a user-chosen baseline cost $c_{\textnormal{base}}$ that encodes prior knowledge or domain-specific structure. This perspective naturally leads to a bilevel optimization problem, where the outer level adjusts the cost parameters and the inner level solves an OT problem whose solution depends on those parameters. Formally, we consider the bilevel problem
$\min\limits_{\theta\in\Theta}\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\varepsilon}(c_{\theta})\big)+\frac{1}{\lambda}\mathscr{D}(c_{\theta},c_{\textnormal{base}})\,\,\textnormal{ s.t. }\,\,\mathbf{\Pi}_{\varepsilon}(c_{\theta})=\operatorname*{arg\,min}\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}_{\theta}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (7)
where $(c_{\theta})_{\theta\in\Theta}$ is a parameterized family of cost functions, $\mathbf{C}_{\theta}:=\big[c_{\theta}(\mathbf{x}_{i},\mathbf{y}_{j})\big]_{ij}$, $c_{\textnormal{base}}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a baseline cost function, and $\mathscr{D}$ is a discrepancy measure between cost functions that encourages the learned cost to remain close to the baseline.
@@
This approach is related to prior work on cost learning in optimal transport [JMLR:v20:18-700], which aims to infer a cost function that explains observed matchings or couplings. Here, however, the objective is different: instead of fitting a cost to reproduce a given transport plan, we learn a cost whose induced OT solution optimizes a downstream criterion, namely fairness.
Problem (7) is well posed because the inner problem is an entropic OT problem, which is strongly convex over the set of couplings $\Pi$ and therefore has a unique solution $\mathbf{\Pi}_{\varepsilon}(c_{\theta})$ for any choice of $\theta$. This uniqueness ensures that the bilevel objective is a well-defined function of the cost parameters. In practice, the bilevel problem can be optimized using gradient-based methods: one can differentiate through the iterative solver (iterative differentiation) [Bolte2023one_step_diff, Franceschi2017ForwardReverse, Maclaurin2015, Pauwels2023], or use (approximate) implicit differentiation to approximate the gradient of the bilevel objective [Dagreou2022SABA, Dagreou2024SRBA, Eisenberger2022, Ghadimi2018].
Our approach can accommodate a wide range of parameterizations for $c_{\theta}$. Two concrete examples are provided below, with the Euclidean distance serving as a natural baseline in both cases.
@@
For a PSD matrix $\mathbf{M}$, the Mahalanobis cost $c_{\mathbf{M}}$ is defined by $c_{\mathbf{M}}(x,y)=(x-y)^{\top}\mathbf{M}(x-y)$. Intuitively, this corresponds to applying a linear transformation to $x$ and $y$ before computing the squared Euclidean distance. A key advantage of this parametrization is its interpretability: the learned matrix $\mathbf{M}$ directly reveals which directions or features in the data are emphasized or downweighted in the transport cost.
@@
A more flexible parametrization uses a neural network $\phi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ to embed the inputs, defining the cost as $c_{\theta}(x,y)=\lVert\phi_{\theta}(x)-\phi_{\theta}(y)\rVert^{2}_{2}$. The network can be trained from scratch or initialized from a pretrained model and fine-tuned for the specific OT task, enabling highly expressive cost functions that capture complex relationships between $x$ and $y$.
Fairness on Unseen Samples.
A key advantage of our cost learning approach is that it produces a reusable cost function, which can be directly applied to match new samples. This naturally raises the question of whether transport plans obtained by solving an optimal transport problem with the learned cost on previously unseen data will maintain a comparable level of fairness.
To address this question, we provide a deviation bound on the fairness of the transport plan that holds uniformly over all cost functions in the parameterized family. Let $\pi_{\varepsilon}^{\star}(c_{\theta})$ denote the optimal transport plan on two measures $\mu$ and $\eta$. Our goal is to control the difference between the fairness of this plan, defined as
$\mathcal{L}_{\mathbf{F}}\big(\pi_{\varepsilon}^{\star}(c_{\theta})\big):=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\bigg(\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi_{\varepsilon}^{\star}(c_{\theta})(x,u,y,t)-\mathbf{F}_{sv}\bigg)^{2}\kern 5.0pt,$
@@
There exists a constant $R_{\Theta}>0$ such that
$\sup\limits_{\theta\in\Theta}\,\,\lVert c_{\theta}\rVert_{\infty}:=\sup\limits_{\theta\in\Theta}\quad\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}c_{\theta}(x,y)<R_{\Theta}.$
This assumption is mildly restrictive. For instance, it holds if $\Theta$ is bounded and the function $(\theta,x,y)\mapsto c_{\theta}(x,y)$ is continuous on $\Theta\times\mathcal{X}\times\mathcal{Y}$. For instance, for the Mahalanobis cost function, the parameter space $\Theta$ is a subset of the space of positive definite matrices, and we have
$R_{\Theta}=\max\limits_{\mathbf{M}\in\Theta}\lVert\mathbf{M}\rVert_{\textnormal{F}}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}$
which is finite under Assumption 3.2 as long $\Theta$ is bounded. Similarly, for the neural cost model, provided that the neural network is $C_{\Theta}$-Lipschitz, we have
$R_{\Theta}=C_{\Theta}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}.$
@@
[Deviation bound for fairness] Assume for simplicity that $n=m$. Under Assumptions 3.2 and 4.2, it holds
$\sup\limits_{\theta\in\Theta}\mathbb{E}\big[\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)-\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\varepsilon^{-1}\big)}{\sqrt{n}}.$
@@
We generalize rigollet2022sample for any bounded cost $c_{\theta}$ by proving that for any test function $\varphi\in L^{\infty}(\mu\otimes\eta)$ and for any $t>0$, we have with probability $1-18e^{-t^{2}}$
$\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}$

### v2

@@ 3.2 Fair Optimal Transport
Under Assumption 3.2, for any target $\mathbf{F}\in\Pi(\mathbf{p},\mathbf{q})$, there exists a unique fair OT plan.
Remark 3.1 (From probabilistic to deterministic fair transport plans).
Given a fair entropic transport plan, a deterministic matching can be generated by sequentially iterating over the rows and, for each row, sampling a column according to the corresponding normalized distribution. For large enough groups, the sampled transport plan will concentrate towards a fair deterministic matching.
Remark 3.2 (Extensions).
Our framework naturally extends to settings in which fairness targets are imposed only on a subset of group pairs, leaving remaining matching probabilities unconstrained, as well as to cost-weighted fairness targets that, for example, balance transport costs across groups.
@@ 3.3 Solving Fair Optimal Transport
where $\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\textstyle\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}$. This result mirrors the foundational insight behind the Sinkhorn algorithm and shows that the fairness constraint can be incorporated via a simple additional projection step. This leads to our modified Sinkhorn algorithm, which we call FairSinkhorn, presented in Algorithm˜1 in which the function $\Phi$ is defined as
$\Phi(\mathbf{u},\mathbf{v})]_{sw}:=\sum_{i\lvert s_{i}=s}\sum_{j\lvert w_{j}=w}\mathbf{u}_{i}\mathbf{v}_{j}\mathbf{K}_{i,j}$
The additional steps relative to the original Sinkhorn algorithm are highlighted in blue.
@@
5:  $\mathbf{u}^{(t+1)}\leftarrow n^{-1}\mathbf{1}\oslash\Big((\mathbf{K}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\odot\mathbf{T}^{(t)}})\mathbf{v}^{(t)}\Big)$
6:  $\mathbf{v}^{(t+1)}\leftarrow m^{-1}\mathbf{1}\oslash\Big((\mathbf{K}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\odot\mathbf{T}^{(t)}})^{\top}\mathbf{u}^{(t+1)}\Big)$
@@
8:  $\mathbf{T}^{(t+1)}\!\leftarrow\!\sum_{sw}\ell^{(t+1)}_{sw}\mathbf{B}_{sw}$ with $\ell^{(t+1)}_{sw}\!=\!\big(\mathbf{L}^{(t+1)}\big)_{sw}$
@@ 4 Two Strategies for Approximately Fair OT
FairSinkhorn enforces exact fairness, which can substantially increase transport cost, sometimes beyond what is acceptable in practice. This “price of fairness” has been studied in operations research for resource allocation problems [bertsimas2011price] and also arises in supervised learning, where enforcing exact group fairness constraints may sacrifice predictive accuracy [zhao2019inherent].
Motivated by these observations, in this section we study approximately fair OT formulations, replacing the linear fairness constraints in (LABEL:eq:pure_empirical_fairness) with a relaxed condition that allows slight deviations from the fairness target while still controlling the total fairness violation:
$\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}):=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\Big(\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\Big)^{2}\leq\rho,$ (4)
where $\rho\geq 0$ is a tolerance level. To find OT plans under this relaxed constraint, we propose two strategies.
@@ 4.1 Fairness-Penalized OT
where $\lambda>0$ controls the strength of the fairness penalty. Importantly, since $\mathcal{L}_{\mathbf{F}}$ is convex, the objective remains strongly convex, ensuring a unique minimizer. This penalized entropic optimal transport problem can be solved efficiently using a generalized conditional gradient algorithm [rakotomamonjy2015generalized]. At each iteration, the algorithm linearizes the fairness regularization around the current iterate and solves a subproblem over the feasible set $\Pi$ to find a search direction. A line search then determines a convex combination of the current iterate and this direction, guaranteeing descent. The subproblem at iteration $t$ corresponds to an entropic OT problem with the modified cost $\mathbf{C}+\nabla\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}^{t})$, which is solved efficiently with the Sinkhorn algorithm. This makes the algorithm computationally practical even for large-scale problems.
Sample complexity of fairness-penalized OT. We establish a sample complexity bound for the above fairness-penalized optimal transport cost by building upon results from entropic optimal transport [genevay2019sample, mena2019statistical]. To formalize this, note that the optimization problem in (5) can be defined for arbitrary measures $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ as (see Section˜C.2 for a formal definition)
$m^{\star}(\mu,\eta)$ $:=\min_{\pi\in\Pi(\mu,\eta)}\textstyle\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)\,d\pi(x,y)$
$\qquad+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)\kern 5.0pt.$
Let $\mu_{n}=\sum_{i=1}^{n}\delta_{x_{i}}$ and $\eta_{n}=\sum_{i=1}^{n}\delta_{y_{i}}$ where $(x_{i})_{1\leq i\leq n}$ and $(y_{i})_{1\leq i\leq n}$ are i.i.d. samples drawn from $\mu$ and $\eta$. In Section˜4.1, we quantify how well we can estimate $m^{\star}(\mu,\eta)$ with $m^{\star}(\mu_{n},\eta_{n})$ as a function of the sample size $n$. Similar to [rigollet2022sample] and [genevay2019sample], our proof relies on the fact that the measures are compactly supported (Assumption 3.2). We also require the following assumption on the ground cost function.
@@
Our result matches the $O(n^{-1/2})$ scaling of sample complexity bounds for standard entropic OT [genevay2019sample, mena2019statistical], showing that adding the fairness penalty does not reduce statistical efficiency. Interesting, the result can be extended to generic convex penalties at the cost of a supplementary $\log(n)$ factor.
@@
Our proof works by finding two adequate random variables $Y_{n}$ and $Z_{n}$ such that
$Y_{n}\!-\!m^{\star}(\mu,\eta)$ $\!\leq\!m^{\star}(\mu_{n},\eta_{n})\!-\!m^{\star}(\mu,\eta)\!\leq\!Z_{n}\!-\!m^{\star}(\mu,\eta).$
Lower bound. We obtain $Y_{n}$ through linearization of the convex penalty, which is guaranteed to yield a proper lower bound through the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized]. This linearized problem can be cast as an entropic OT problem with modified cost, allowing us to leverage existing sample complexity results [genevay2019entropy, mena2019statistical].
Upper bound. To obtain $Z_{n}$ we evaluate the fairness-penalized loss at the minimizer $\hat{\pi}_{n}^{\star}$ of the linearized problem, that is, we consider
$m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)$
$\hskip-100.00015pt\leq\!\langle c,\hat{\pi}_{n}^{\star}\rangle\!+\!\varepsilon\textbf{KL}(\hat{\pi}_{n}^{\star}|\mu_{n}\!\otimes\!\eta_{n})\!+\!\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})\!-\!m^{\star}(\mu,\eta),$ (6)
where the upper bound follows from the fact that $\hat{\pi}_{n}^{\star}$ satisfies the constraints defining $m^{\star}(\mu_{n},\eta_{n})$. To establish that the expectation of the absolute value of (4.1) vanishes at the rate of $1/\sqrt{n}$, we again make use of the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized] to reduce the problem to the sample complexity of entropic OT. We handle the part $\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta)$ through a generalization of Theorem 6 from [rigollet2022sample]. ∎
@@ 4.2 Fair OT via Cost Learning
The key idea behind our second strategy is to learn a ground cost function $c_{\theta}$ so that the entropic OT solution $\mathbf{\Pi}(c_{\theta})$ it induces minimizes the fairness violation. In other words, rather than modifying the transport plan directly, we reshape the geometry of the OT problem. This perspective leads naturally to a bilevel optimization problem, with an outer level that adjusts the cost parameters and an inner level that solves an OT problem whose solution depends on those parameters. Formally, we consider
$\begin{aligned} &\min\limits_{\theta\in\Theta}\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\varepsilon}(c_{\theta})\big)+\frac{1}{\lambda}\mathscr{D}(c_{\theta},c_{\textnormal{base}})\\ &\textnormal{s.t.}\,\,\mathbf{\Pi}_{\varepsilon}(c_{\theta})=\operatorname*{arg\,min}\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}_{\theta}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,\end{aligned}$ (7)
where $(c_{\theta})_{\theta\in\Theta}$ is a parameterized family of cost functions, $\mathbf{C}_{\theta}:=\big[c_{\theta}(\mathbf{x}_{i},\mathbf{y}_{j})\big]_{ij}$, $c_{\textnormal{base}}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a baseline cost encoding prior knowledge or domain-specific structure, and $\mathscr{D}$ is a discrepancy measure that encourages the learned cost to remain close to the baseline, thereby controlling deviations from the original OT problem.
@@
This formulation is related to prior work on cost learning in optimal transport [JMLR:v20:18-700, andrade2023sparsistency], which aims to infer a cost function that explains observed matchings or couplings. Here, however, the objective is different: instead of fitting a cost to reproduce a given transport plan, we learn a cost whose induced OT solution optimizes a downstream criterion, namely fairness.
Problem (7) is well posed since the inner entropic OT problem is strongly convex over $\Pi$, yielding a unique solution $\mathbf{\Pi}_{\varepsilon}(c_{\theta})$ for any $\theta$ and thus a well-defined bilevel objective. This bilevel problem can be optimized using gradient-based methods by differentiating through an iterative solver (iterative differentiation) [Bolte2023one_step_diff, Franceschi2017ForwardReverse, Maclaurin2015, Pauwels2023] or using implicit differentiation to approximate the gradient of the bilevel objective [Dagreou2022SABA, Dagreou2024SRBA, Eisenberger2022, Ghadimi2018].
Our approach can accommodate a wide range of parameterizations for $c_{\theta}$.
@@
Assume $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{d}$. For a PSD matrix $\mathbf{M}$, the Mahalanobis cost $c_{\mathbf{M}}$ is defined by $c_{\mathbf{M}}(x,y)=(x-y)^{\top}\mathbf{M}(x-y)$. This corresponds to applying a linear transformation to $x$ and $y$ before computing the squared Euclidean distance. A key advantage of this parametrization is its interpretability: the learned matrix $\mathbf{M}$ directly reveals which directions or features in the data are emphasized or downweighted in the transport cost.
@@
A more flexible parametrization uses neural networks $\phi_{\theta_{1}}:\mathcal{X}\to\mathbb{R}^{k}$ and $\phi_{\theta_{2}}:\mathcal{Y}\to\mathbb{R}^{k}$ to embed the inputs, defining the cost as $c_{\theta}(x,y)=\lVert\phi_{\theta_{1}}(x)-\phi_{\theta_{2}}(y)\rVert^{2}_{2}$. The networks can be trained from scratch or initialized from a pretrained model and fine-tuned for the specific OT task, enabling highly expressive cost functions that capture complex relationships between $x$ and $y$.
Fairness on unseen samples. A key advantage of our cost learning approach is that it produces a reusable cost function, which can be directly applied to match new samples. This naturally raises the question of whether transport plans obtained by solving an optimal transport problem with the learned cost on previously unseen data will maintain a comparable level of fairness. To address this question, we derive a pointwise deviation bound on the fairness of the transport plan computed from finite samples relative to the population solution, uniformly over the parameterized cost family. Let $\pi_{\varepsilon}^{\star}(c_{\theta})$ denote the OT plan on two measures $\mu$ and $\eta$. Our goal is to control the difference between the fairness of this plan, defined as
$\mathcal{L}_{\mathbf{F}}\big(\pi_{\varepsilon}^{\star}(c_{\theta})\big)$
$\hskip-100.00015pt:=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}$ $\Big(\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi_{\varepsilon}^{\star}(c_{\theta})(x,u,y,t)-\mathbf{F}_{sv}\Big)^{2},$
@@
There exists a constant $R_{\Theta}>0$ s.t.
$\sup\limits_{\theta\in\Theta}\,\,\lVert c_{\theta}\rVert_{\infty}:=\sup\limits_{\theta\in\Theta}~~\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}c_{\theta}(x,y)<R_{\Theta}.$
This assumption is mildly restrictive. For instance, it holds if $\Theta$ is bounded and the function $(\theta,x,y)\mapsto c_{\theta}(x,y)$ is continuous on $\Theta\times\mathcal{X}\times\mathcal{Y}$. For instance, for the Mahalanobis cost function (Example 4.1), the parameter space $\Theta$ is a subset of the space of positive definite matrices and we have
$R_{\Theta}=\max\limits_{\mathbf{M}\in\Theta}\lVert\mathbf{M}\rVert_{\textnormal{F}}\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}\lVert x-y\rVert_{2}$
which is finite under Assumption 3.2 as long $\Theta$ is bounded. Similarly, for the neural cost function (Example 4.2), provided that the neural networks are $C_{\Theta}$-Lipschitz, we have
$R_{\Theta}=C_{\Theta}\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}\lVert x-y\rVert_{2}\kern 5.0pt.$
@@
Assume for simplicity that $n=m$. Under Assumptions 3.2 and 4.2, we have
$\sup\limits_{\theta\in\Theta}\mathbb{E}\big[\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)\!-\!\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\big]\!\lesssim\!\frac{\exp\big(\frac{5R_{\Theta}}{\varepsilon}\big)}{\sqrt{n}}\kern 5.0pt.$
@@
We generalize rigollet2022sample to any bounded cost $c_{\theta}$ by proving that for any test function $\varphi\in L^{\infty}(\mu\otimes\eta)$ and for any $t>0$, we have with probability $1-18e^{-t^{2}}$
$\big\lvert\!\textstyle\int\!\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\lesssim\displaystyle\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}$